# All in One: Visual-Description-Guided Unified Point Cloud Segmentation

### v1

@@ 1 Introduction
Figure 1: Comparison between existing multi-modal 3D segmentation methods and our proposed approach. (a) Existing methods depend on paired images and captioning models to generate descriptions, requiring real-time 2D-3D alignment, which increases computational overhead. (b) Our approach leverages pre-generated class descriptions from LLMs and reference images retrieved from the internet, establishing static semantic anchors that enhance efficiency and robustness.
@@
In particular, unlike open-vocabulary methods [14, 57] that require real-time fusion of 3D scenes with paired images and text captions, our approach follows a closed-set paradigm with predefined class knowledge. We establish static semantic anchors through offline-generated class descriptions and reference images (Fig. 1). During inference, our method operates without paired images, captioning models, or additional vision-language modules, enhancing efficiency and deployment practicality. By integrating static multimodal knowledge with spatially aware feature learning, our approach achieves superior class separation and instance delineation compared to conventional closed-set 3D segmentation baselines. Both quantitative and qualitative comparisons demonstrate consistent improvements in our model performance in 3D instance, semantic, and panoptic segmentation tasks.
@@ 3.2 Multi-Modal Reference Queries
$\mathbf{D}^{c}=[d^{c}_{1},\dots,d^{c}_{K}]=\text{LLM-DesGen}(c),$ (3)
where $\mathbf{D}^{c}$ represents the descriptions of the $c$-th class, and $d^{c}_{k}$ denotes the $k$-th description of the $c$-th class. Finally, we can get the set of descriptions for all classes: $\mathbf{D}=[\mathbf{D}^{1},\dots,\mathbf{D}^{C}]$. Next, we utilize a text encoder (i.e., CLIP [40]) to embed these descriptions into description embeddings:
$\mathbf{Q}_{d}=\text{CLIP-T}(\mathbf{D}),$ (4)
where $\mathbf{Q}_{d}\in\mathbb{R}^{C\times K\times d}$ is the description embedding of all classes.
@@ 3.3 Multi-Modal Unified Mask Decoder
$\!\!\!\mathbf{M}_{ins}=D(\mathbf{Q}_{ins},\widetilde{\mathbf{X}}),\mathbf{M}_{d}=D(\mathbf{Q}_{d},\widetilde{\mathbf{X}}),\mathbf{M}_{o}=D(\mathbf{Q}_{o},\widetilde{\mathbf{X}}),\!\!\!$ (6)
where $\mathbf{M}_{ins}\in\mathbb{R}^{N\times S}$ is the predicted instance segmentation mask, $\mathbf{M}_{d}\in\mathbb{R}^{N\times C\times K}$ and $\mathbf{M}_{o}\in\mathbb{R}^{N\times C\times L}$ are the semantic segmentation masks predicted by the description and image queries, respectively.
@@
$\widehat{\mathbf{M}}_{d}^{m}=\max_{k}(\mathbf{M}_{d}^{m}).$ (7)
@@
${\mathbf{M}}_{sem}^{m}=\widehat{\mathbf{M}}_{d}^{m}+\widehat{\mathbf{M}}_{o}^{m},$ (8)
@@ 3.4 Training Losses
$\mathcal{L}_{\text{SVC}}=\mathcal{L}_{\text{c}}(\widetilde{\mathbf{X}},\mathbf{Q}_{d})+\mathcal{L}_{\text{c}}(\widetilde{\mathbf{X}},\mathbf{Q}_{o}),$ (9)
@@
where $\mathbf{q}_{+}$ denotes the positive description or image embedding corresponding to the same class as the point feature $\tilde{\mathbf{x}}_{i}$, $\operatorname{sim}(\cdot)$ denotes the inner product similarity, and $\tau$ is the temperature parameter. This loss aligns point features with class-specific embeddings, enhancing both semantic consistency and class discriminability. It also facilitates better integration of information from reference descriptions and images, contributing to a more robust segmentation.
The overall loss function for the unified 3D segmentation is defined as:
@@ 4.2 Comparisons with SOTA Methods
We compare our method against state-of-the-art (SOTA) approaches on three datasets. Table 1 presents comparisons on the S3DIS [2] dataset. Across both the Area-5 and 6-fold cross-validation benchmarks, our method demonstrates strong competitive performance compared to SOTA methods. Notably, on the Area-5 benchmark, our method achieves significant improvements in instance segmentation, outperforming previous methods by a large margin. Specifically, VDG-Uni3DSeg achieves a 1.9-point gain in mAP50 and a 3.0-point improvement in mPrec50 over the best-performing baseline. For panoptic segmentation, our method surpasses SOTA methods with a 3.1-point gain in PQ and a remarkable 7.0-point improvement in PQth. Additionally, for semantic segmentation, VDG-Uni3DSeg achieves a 1.0-point improvement in mIoU over our baseline, OneFormer3D [27]. On the 6-fold cross-validation benchmark, VDG-Uni3DSeg achieves similarly notable improvements across all metrics, demonstrating its robustness and effectiveness in different settings.
@@
Table 2 presents comparisons on the ScanNet [13] dataset, including results for both the validation split and the hidden test split. In the validation split, our method consistently improves performance across all segmentation tasks, including instance, semantic, and panoptic segmentation. Specifically, for instance segmentation, VDG-Uni3DSeg achieves a 0.9-point improvement in mAP50, demonstrating its enhanced ability to detect and distinguish individual objects. For semantic segmentation, our method achieves a 0.6-point improvement in mIoU, highlighting its improved capability to classify points accurately across the scene. Furthermore, for panoptic segmentation, VDG-Uni3DSeg achieves a 0.8-point improvement in PQth, demonstrating its effectiveness in jointly optimizing instance and semantic predictions. Additionally, VDG-Uni3DSeg also achieves improvements in the ScanNet hidden test set, further highlighting its strong generalization ability. Table 3 presents the results for ScanNet200 [42], which introduces fine-grained classes. VDG-Uni3DSeg achieves state-of-the-art performance in instance and panoptic segmentation, demonstrating its robustness in complex segmentation tasks. Although our model achieves slightly lower semantic mIoU, it remains competitive and maintains balanced performance across all tasks.
@@ 4.3 Ablation Study
Table 4 presents an ablation study evaluating the contribution of each component in our VDG-Uni3DSeg framework on S3DIS. First, removing the Spatial Enhancement Module slightly degrades the performance, particularly in instance and panoptic segmentation metrics (e.g., a 2-point decrease in mAP and a 0.7-point drop in PQ). This underscores the importance of spatial information in effectively distinguishing instances within the scene. Next, we analyze the impact of reference queries. Removing either the Description Query or the Image Query leads to a significant performance drop, with a more pronounced decline when the Description Query is removed. This demonstrates the benefit of detailed descriptions generated by LLMs in our method. Furthermore, incorporating reference images helps bridge the gap between textual descriptions and the 3D point cloud. By integrating description and image queries, our method enables each point to associate with the most relevant class representation, leading to overall improvements across all segmentation tasks. Finally, removing the SVC Loss causes the most significant drop in instance and panoptic segmentation metrics, including a 2.3-point decline in mAP25 and a 4.2-point drop in PQth. This demonstrates that SVC Loss encourage the model to learn more discriminative segmentation features by leveraging multi-modal queries as anchors for feature alignment. In general, each component of our method contributes to improving feature alignment and overall segmentation performance.
@@ Appendix A Ablation Study
128 78.1 58.8 86.1 78.0 76.3 71.6
Table R5: Effect of different sample size on ScanNet.
@@
Ours 73.9 60.1 82.7 73.9 70.8 65.3
Table R6: Comparision with random images on S3DIS.

### v2

@@ 1 Introduction
Figure 1: Comparison between existing multi-modal 3D segmentation methods and our proposed approach. (a) Existing methods depend on paired images and captioning models to generate descriptions, requiring real-time 2D-3D alignment, which increases computational overhead. (b) Our approach leverages pre-generated class descriptions from LLMs and reference images retrieved from the internet, establishing offline semantic references that enhance efficiency and robustness.
@@
In particular, unlike open-vocabulary methods [14, 57] that require real-time fusion of 3D scenes with paired images and text captions, our approach follows a closed-set paradigm with predefined class knowledge. We establish offline semantic references through offline-generated class descriptions and reference images (Fig. 1). During inference, our method operates without paired images, captioning models, or additional vision-language modules, enhancing efficiency and deployment practicality. By integrating offline multimodal knowledge with spatially aware feature learning, our approach achieves superior class separation and instance delineation compared to conventional closed-set 3D segmentation baselines. Both quantitative and qualitative comparisons demonstrate consistent improvements in our model performance in 3D instance, semantic, and panoptic segmentation tasks.
@@ 3.2 Multi-Modal Reference Queries
$\mathbf{T}^{c}=[t^{c}_{1},\dots,t^{c}_{K}]=\text{LLM-DesGen}(c),$ (3)
where $\mathbf{T}^{c}$ represents the descriptions of the $c$-th class, and $t^{c}_{k}$ denotes the $k$-th description of the $c$-th class. Finally, we can get the set of descriptions for all classes: $\mathbf{T}=[\mathbf{T}^{1},\dots,\mathbf{T}^{C}]$. Next, we utilize a text encoder (i.e., CLIP [40]) to embed these descriptions into description embeddings:
$\mathbf{Q}_{t}=\text{CLIP-T}(\mathbf{T}),$ (4)
where $\mathbf{Q}_{t}=[\mathbf{Q}_{t}^{1},\dots,\mathbf{Q}_{t}^{C}]$, where $\mathbf{Q}_{t}^{c}\in\mathbb{R}^{K\times d}$ denotes the $K$ queries for class $c$
@@ 3.3 Multi-Modal Unified Mask Decoder
$\!\!\!\mathbf{M}_{ins}=D(\mathbf{Q}_{ins},\widetilde{\mathbf{X}}),\mathbf{M}_{t}=D(\mathbf{Q}_{t},\widetilde{\mathbf{X}}),\mathbf{M}_{o}=D(\mathbf{Q}_{o},\widetilde{\mathbf{X}}),\!\!\!$ (6)
where $\mathbf{M}_{ins}\in\mathbb{R}^{N\times S}$ is the predicted instance segmentation mask, $\mathbf{M}_{t}\in\mathbb{R}^{N\times C\times K}$ and $\mathbf{M}_{o}\in\mathbb{R}^{N\times C\times L}$ are the semantic segmentation masks predicted by the description and image queries, respectively.
@@
$\widehat{\mathbf{M}}_{t}^{m}=\max_{k}(\mathbf{M}_{t}^{m}).$ (7)
@@
${\mathbf{M}}_{sem}^{m}=\widehat{\mathbf{M}}_{t}^{m}+\widehat{\mathbf{M}}_{o}^{m},$ (8)
@@ 3.4 Training Losses
$\mathcal{L}_{\text{SVC}}=\mathcal{L}_{\text{c}}(\widetilde{\mathbf{X}},\mathbf{Q}_{t})+\mathcal{L}_{\text{c}}(\widetilde{\mathbf{X}},\mathbf{Q}_{o}),$ (9)
@@
where $\mathbf{q}_{+}$ denotes the positive description or image embedding corresponding to the same class as the point feature $\tilde{\mathbf{x}}_{i}$, $\operatorname{sim}(\cdot)$ denotes the inner product similarity, and $\tau$ is the temperature parameter, empirically set to 1.0. This loss aligns point features with class-specific embeddings, enhancing both semantic consistency and class discriminability. It also facilitates better integration of information from reference descriptions and images, contributing to a more robust segmentation.
The overall loss function for our method is defined as:
@@ 4.2 Comparisons with SOTA Methods
We compare our method against state-of-the-art (SOTA) approaches on three datasets. Table 1 presents comparisons on the S3DIS [2] dataset. Across both the Area-5 and 6-fold cross-validation benchmarks, our method demonstrates strong competitive performance compared to SOTA methods. Notably, on the Area-5 benchmark, our method achieves significant improvements in instance segmentation, outperforming previous methods by a large margin. Specifically, VDG-Uni3DSeg achieves a 2.1-point gain in mAP50 and a 1.4-point improvement in mAP over the best-performing baseline. For panoptic segmentation, our method surpasses SOTA methods with a 4.1-point gain in PQ and a remarkable 9.6-point improvement in PQth. Additionally, for semantic segmentation, VDG-Uni3DSeg achieves a 1.7-point improvement in mIoU over our baseline, OneFormer3D [27]. On the 6-fold cross-validation benchmark, VDG-Uni3DSeg achieves similarly notable improvements across all metrics, demonstrating its robustness and effectiveness in different settings.
@@
Table 2 presents comparisons on the ScanNet [13] dataset, including results for both the validation split and the hidden test split. In the validation split, our method consistently improves performance across all segmentation tasks, including instance, semantic, and panoptic segmentation. Specifically, for instance segmentation, VDG-Uni3DSeg achieves a 1.3-point improvement in mAP50, demonstrating its enhanced ability to detect and distinguish individual objects. For semantic segmentation, our method achieves a 0.4-point improvement in mIoU, highlighting its improved capability to classify points accurately across the scene. Furthermore, for panoptic segmentation, VDG-Uni3DSeg also achieves improvement in PQ, demonstrating its effectiveness in jointly optimizing instance and semantic predictions. Additionally, VDG-Uni3DSeg also achieves improvements in the ScanNet hidden test set, further highlighting its strong generalization ability. Table 3 presents the results for ScanNet200 [42], which introduces fine-grained classes. VDG-Uni3DSeg achieves state-of-the-art performance in all instance, semantic, and panoptic segmentation, demonstrating its robustness in complex segmentation tasks. Although our model achieves slightly lower semantic mIoU, it remains competitive and maintains balanced performance across all tasks.
@@ 4.3 Ablation Study
Table 4 presents an ablation study evaluating the contribution of each component in our VDG-Uni3DSeg framework on S3DIS. First, removing the Spatial Enhancement Module slightly degrades the performance, particularly in instance and panoptic segmentation metrics (e.g., a 2-point decrease in mAP and a 1.7-point drop in PQ). This underscores the importance of spatial information in effectively distinguishing instances within the scene. Next, we analyze the impact of reference queries. Removing either the Description Query or the Image Query leads to a significant performance drop, with a more pronounced decline when the Description Query is removed. This demonstrates the benefit of detailed descriptions generated by LLMs in our method. Furthermore, incorporating reference images helps bridge the gap between textual descriptions and the 3D point cloud. By integrating description and image queries, our method enables each point to associate with the most relevant class representation, leading to overall improvements across all segmentation tasks. Finally, removing the SVC Loss causes the most significant drop in instance and panoptic segmentation metrics, including a 2.6-point decline in mAP25 and a 6.9-point drop in PQth. This demonstrates that SVC Loss encourage the model to learn more discriminative segmentation features by leveraging multi-modal queries as anchors for feature alignment. In general, each component of our method contributes to improving feature alignment and overall segmentation performance.
@@ Appendix A Ablation Study
Table R5: Effect of different sample size on ScanNet.
@@
128 78.5 59.3 85.7 78.3 76.2 71.5
Table R6: Comparision with random images on S3DIS.
@@
Ours 74.1 60.1 81.0 73.9 71.5 66.3